# DexDiffuser: Generating Dexterous Grasps with Diffusion Models

### v2

@@ Abstract
We introduce DexDiffuser, a novel dexterous grasping method that generates, evaluates, and refines grasps on partial object point clouds. DexDiffuser includes the conditional diffusion-based grasp sampler DexSampler and the dexterous grasp evaluator DexEvaluator. DexSampler generates high-quality grasps conditioned on object point clouds by iterative denoising of randomly sampled grasps. We also introduce two grasp refinement strategies: Evaluator-Guided Diffusion and Evaluator-based Sampling Refinement. The experiment results demonstrate that DexDiffuser consistently outperforms the state-of-the-art multi-finger grasp generation method FFHNet with an, on average, 9.12% and 19.44% higher grasp success rate in simulation and real robot experiments, respectively. Supplementary materials are available at https://yulihn.github.io/DexDiffuser_page/
@@ I Introduction
In this work, we propose the new data-driven dexterous grasping method DexDiffuser to address the dimensionality issue. DexDiffuser includes DexSampler, a new conditional diffusion-based dexterous grasp sampler, and DexEvaluator, a new dexterous grasp evaluator. DexSampler generates high-quality grasps conditioned on object point clouds by iteratively denoising randomly sampled grasps. In addition to DexDiffuser, we also propose two grasp refinement strategies: Evaluator-Guided Diffusion (EGD), which is only applicable for diffusion models, and Evaluator-based Sampling Refinement (ESR), which applies to all dexterous grasp sampling methods.
We experimentally assess DexDiffuser’s capability to sample, evaluate, and refine 16- Degrees of Freedom (DoF) Allegro Hand grasps on three visually and geometrically distinctive object datasets in simulation and on one dataset in the real world. DexDiffuser is benchmarked against the state-of-the-art multi-finger grasp generation method FFHNet [2]. The experimental results indicate that our best method achieves a grasp success rate of 98.77% in simulation and 68.89% in the real world, which is 11.63% and 20.00% higher than the respectively best FFHNet model.
@@
  • A dataset of 1.7 million successful and unsuccessful Allegro Hand grasp across 5378 objects (Section V).
@@ II-A Data-Driven Dexterous Grasping
Data-driven dexterous grasp sampling methods primarily fall into three categories: 1) those that generate a contact map on the object’s surface [5, 6, 7], 2) those that rely on shape completion [8, 9, 10, 11, 12, 13, 14, 15], or 3) those that train a grasping policy based on Reinforcement Learning (RL) [16, 17] or human demonstrations [18]. The main limitation of the methods in 1) is the assumption of complete object observation, which hinders real-world applications where complete visibility cannot be guaranteed. Methods in 2) are limited because the generated grasp quality is closely intertwined with the completion quality while collecting expensive demonstrations, and the sim-to-real gap limits the methods in 3).
@@ IV-A The Basis Point Set Representation
DexDiffuser generates and samples grasps directly on partial object point clouds. In this work, $\mathbf{O}$ is encoded into $\mathbf{f}_{\mathbf{O}}$ using Basis Point Set (BPS) [30] as this encoding has shown to work well in prior dexterous grasping work [2]. The BPS encoding represents the shortest distance between each point in a fixed set of basis points and all points in $\mathbf{O}$. The benefit of BPS is that it captures the shape and spatial properties of the object and is of fixed length, which allows for the use of computationally efficient Fully Connected (FC) layers to further process $\mathbf{f_{O}}$.
@@ IV-B Diffusion-based Grasp Sampler
where $\hat{\mathbf{\epsilon}}{{}_{t}}\sim\mathcal{N}(\mathbf{g}{{}_{t-1}};\hat{\mu}_{\bm{\theta}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}}),\hat{\Sigma}_{{\bm{\theta}}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}}))$ and $\mathbf{\epsilon}_{t}$ is the ground-truth noise.
@@ IV-D1 EGD
$\mathbf{g}{{}_{\text{guide}}}=$ $\nabla_{\mathbf{g}{{}_{t}}}\log\mathcal{D}_{\phi}(\hat{y}=1|\mathbf{g}{{}_{t}},\mathbf{f_{O}})+\log(T-t+1),$ (9)
@@ V Dataset
The final dataset consisted of 1.7 million grasps, where approximately 40.52% were successful and 59.48% were unsuccessful. We split this dataset into a training set of 4303 objects and a test set of 1,075. We further add 49 objects from the EGAD! dataset [40] and 16 objects from the MultiDex dataset [5] to the the test dataset.
@@ VI Experimental evaluation
We experimentally evaluated DexDiffuser in both simulation and the real world, comparing it to the state-of-the-art grasp sampler and evaluator FFHNet [2]. The specific questions we wanted to address with the experiments were:
@@
    What effect does the BPS-encoding have on generating successful grasps?
@@ VI-A Simulation Experiments
All simulation experiments were conducted in Isaac Gym [41] following the setup from [5]. Specifically, test objects were initialized as free-floating objects, and a virtual 16-DoF Allegro Hand was used for grasping each of them. For each of the 1140 test objects, 10 point clouds were rendered from randomly sampled camera poses. Then, all methods generated 20 grasps per point cloud, which amounted to 228,000 test grasps in total. Once an object was grasped, the grasp stability was assessed by applying external forces along the six directions in space for 50 timesteps. A grasp succeeded if the object remained in the hand throughout the stability test.
@@
Figure 2: Generated grasp examples. Qualitative evaluation of DexSampler-BPS on objects from different datasets. The grasps, shown in pink, are generated on the partially observed point clouds, shown in red, rendered from the complete meshes, shown in blue.
@@
{adjustbox}
max width= Method Dataset MultiDex(%$\uparrow$ EGAD! (%$\uparrow$ DexGN (%$\uparrow$ FFHNet [2] 72.47 87.14 82.16 DexS-PN2 82.28 98.16 81.77 DexS-BPS 87.28 98.62 82.64 FFHNet-ESR-1 71.91 86.77 82.05 FFHNet-ESR-2 72.16 86.86 82.21 DexS-BPS-ESR-1 87.16 98.67 82.57 DexS-BPS-ESR-2 87.47 98.68 82.60 DexS-BPS-EGD 87.63 98.68 82.88 DexS-BPS-EGD-ESR-1 87.41 98.68 82.94 DexS-BPS-EGD-ESR-2 87.78 98.77 82.90
TABLE II: Grasp success rate of different methods. The three first rows exclude grasp refinement, while the bottom nine include it. DexS represents the DexSampler. DexGN is short for the DexGraspNet test split. $\uparrow$: The higher, the better.
{adjustbox}
TABLE III: Experimental results for real-world grasping. DexS represents DexSampler.
@@ VI-A2 Grasp Sampling
To address the second question, we evaluated both FFHNet and DexSampler using either the PointNet++ or BPS encoding. The simulation results are presented in Table II. These results demonstrate that DexSampler, irrespective of the encoding method, consistently outperformed the FFHNet generator across all test datasets. The superior performance underscores the diffusion models’ effectiveness in synthesizing successful grasps. Moreover, DexSampler-BPS performed much better than DexSampler-PN2, indicating that the BPS encoding is less affected by point cloud irregularities, including missing parts and non-uniform density distribution. Finally, Fig. 2 shows some grasps generated by the DexSampler-BPS on the three datasets.
@@ VI-B Real Robot Experiment
Figure 3: Experimental objects. From 1 to 9: cracker box, sugar box, mustard bottle, bleach cleanser, sprayer, metal mug, goblet, toy plane, and pringles
Figure 4: Successful and failed grasp examples from DexSampler-BPS-ESR-2. The first two images successfully grasp the bleach cleanser and the goblet, while the latter two show different failure cases on the metal mug and the pringles.
We evaluated the following methods: FFHNet, FFHNet-ESR-2, DexDiffuser-BPS, DexDiffuser-BPS-ESR-2, DexDiffuser-BPS-EGD and DexDiffuser-BPS-EGD-ESR-2. It is worth highlighting that these methods are only trained on simulated data. The experimental results are summarized in Table III. These results, like the simulation results, show that DexDiffuser consistently outperformed FFHNet across all test objects. The best method for picking real-world objects was DexSampler-BPS-ESR-2, and some of the grasps it produced are shown in Fig. 4. Interestingly, no methods successfully grasped the toy plane, mainly because the Allegro hand was not strong enough to keep the object secure when the last joint rotated $\pm$90°, which is demonstrated in the accompanying video [^1] https://yulihn.github.io/DexDiffuser_page/.
@@
We also report the inference time in Table III. The iterative denoising process makes the diffusion-based methods slower at inference than FFHNet. For instance, the best performing DexDiffuser (DexS-BPS-ESR-2) is 3.5 times slower than the best performing FFHNet (FFHNet-ESR-2).

### v3

@@ Abstract
We introduce DexDiffuser, a novel dexterous grasping method that generates, evaluates, and refines grasps on partial object point clouds. DexDiffuser includes the conditional diffusion-based grasp sampler DexSampler and the dexterous grasp evaluator DexEvaluator. DexSampler generates high-quality grasps conditioned on object point clouds by iterative denoising of randomly sampled grasps. We also introduce two grasp refinement strategies: Evaluator-Guided Diffusion and Evaluator-based Sampling Refinement. The experiment results demonstrate that DexDiffuser consistently outperforms the state-of-the-art multi-finger grasp generation method FFHNet with an, on average, 9.12% and 19.44% higher grasp success rate in simulation and real robot experiments, respectively. Supplementary materials are available at https://yulihn.github.io/DexDiffuser_page/.
@@ I Introduction
In this work, we propose the new data-driven dexterous grasping method DexDiffuser to address the dimensionality issue. DexDiffuser includes DexSampler, a new conditional diffusion-based dexterous grasp sampler, and DexEvaluator, a new dexterous grasp evaluator. DexSampler generates high-quality grasps conditioned on object point clouds by iteratively denoising randomly sampled grasps. We trained the models on 1.7 million successful and unsuccessful Allegro Hand grasp generated using DexGraspNet [4] across 5378 objects. In addition to DexDiffuser, we also propose two grasp refinement strategies: Evaluator-Guided Diffusion (EGD), which is only applicable for diffusion models, and Evaluator-based Sampling Refinement (ESR), which applies to all dexterous grasp sampling methods.
We experimentally assess DexDiffuser’s capability to sample, evaluate, and refine 16- Degrees of Freedom (DoF) Allegro Hand grasps on three visually and geometrically distinctive object datasets in simulation and on one dataset in the real world. DexDiffuser is benchmarked against the state-of-the-art contactmap-based method UniDexGrasp [5] in simulation and the multi-finger grasp generation method FFHNet [2] in simulation and the real world. The experimental results indicate that our best method achieves a grasp success rate of 98.77% in simulation and 68.89% in the real world, which is 11.63% and 20.00% higher than the respectively best FFHNet model.
@@ II-A Data-Driven Dexterous Grasping
Data-driven dexterous grasp sampling methods primarily fall into three categories: 1) those that generate a contact map on the object’s surface [5, 7, 8, 9], 2) those that rely on shape completion [10, 11, 12, 13, 14, 15, 16, 17], or 3) those that train a grasping policy based on Reinforcement Learning (RL) [18, 19] or human demonstrations [20]. The main limitation of the methods in 1) is the assumption of complete object observation, which hinders real-world applications where complete visibility cannot be guaranteed. Methods in 2) are limited because the generated grasp quality is closely intertwined with the completion quality while collecting expensive demonstrations, and the sim-to-real gap limits the methods in 3).
@@ IV-A The Basis Point Set Representation
DexDiffuser generates and samples grasps directly on partial object point clouds $\mathbf{O}$. In this work, $\mathbf{O}$ is encoded into $\mathbf{f}_{\mathbf{O}}$ using Basis Point Set (BPS) [32] as this encoding has shown to work well in prior dexterous grasping work [2]. The BPS encoding represents the shortest distance between each point in a fixed set of basis points and all points in $\mathbf{O}$. The benefit of BPS is that it captures the shape and spatial properties of the object and is of fixed length, which allows for the use of computationally efficient Fully Connected (FC) layers to further process $\mathbf{f_{O}}$.
@@ IV-B Diffusion-based Grasp Sampler
where $\hat{\mathbf{\epsilon}}{{}_{t}}=\epsilon_{\theta}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}})$ and $\mathbf{\epsilon}_{t}$ is the ground-truth noise.
@@ IV-D1 EGD
$\mathbf{g}{{}_{\text{guide}}}=$ $\log(T-t+1)\nabla_{\mathbf{g}{{}_{t}}}\log\mathcal{D}_{\phi}(\hat{y}=1|\mathbf{g}{{}_{t}},\mathbf{f_{O}}),$ (9)
@@ V Dataset
The final dataset consisted of 1.7 million grasps, where approximately 40.52% were successful and 59.48% were unsuccessful. We split this dataset into a training set of 4303 objects and a test set of 1,075. We further add 49 objects from the EGAD! dataset [41], 16 objects from the MultiDex dataset [7], and 12 objects from the KIT dataset [42] to the the test dataset.
@@ VI Experimental evaluation
We experimentally evaluated DexDiffuser in both simulation and the real world. The specific questions we wanted to address with the experiments were:
@@
    How well can DexSampler generate successful and diverse grasps?
@@ VI-A Simulation Experiments
All simulation experiments were conducted in Isaacgym [43] following the setup from [7] [^1] It is worth pointing out that the criteria is less strict than [4].. Specifically, test objects were initialized as free-floating objects, and a virtual 16-DoF Allegro Hand was used for grasping each of them. For each of the 1152 test objects, 10 point clouds were rendered from randomly sampled camera poses. Then, all methods generated 20 grasps per point cloud, which amounted to 230,400 test grasps in total. Once an object was grasped, the grasp stability was assessed by applying external forces along the six directions in space for 50 timesteps. A grasp succeeded if the object remained in the hand throughout the stability test.
@@ VI-A2 Grasp Sampling
{adjustbox}
max width= Method Dataset MultiDex (%$\uparrow$ EGAD! (%$\uparrow$ DexGN-D (%$\uparrow$ KIT (%$\uparrow$ FFHNet [2] 72.47 87.14 82.16 68.04 DexS-PN2 82.28 98.16 81.77 81.13 DexS-BPS 87.28 98.62 82.64 83.13 UniDexGrasp w/o TTA [5] 71.25 89.23 54.19 68.38 FFHNet-ESR-1 71.91 86.77 82.05 67.71 FFHNet-ESR-2 72.16 86.86 82.21 68.29 UniDexGrasp 72.47 85.35 61.70 67.46 DexS-BPS-ESR-1 87.16 98.67 82.57 83.58 DexS-BPS-ESR-2 87.47 98.68 82.60 83.50 DexS-BPS-EGD 87.63 98.68 82.88 82.96 DexS-BPS-EGD-ESR-1 87.41 98.68 82.94 83.08 DexS-BPS-EGD-ESR-2 87.78 98.77 82.90 82.96
TABLE II: Grasp success rate of different methods. The four first rows exclude grasp refinement, while the bottom eight include it. DexS represents the DexSampler. DexGN-D is short for the DexGraspNet dataset test split. For UniDexGrasp, w/o TTA means without the test-time contact-based optimization. $\uparrow$: The higher, the better.
{adjustbox}
[…]
TABLE III: Grasp diversity and max penetration depth. For diversity, we average the diversity of all datasets. DexGN-S represents DexGraspNet sampler. $\uparrow$: The higher, the better. $\downarrow$: The lower, the better.
To address the second question, we compared DexSampler using the PointNet++ and the BPS encoding against FFHNet and the grasp proposal network from UniDexGrasp [5]. It is worth pointing out that UniDexGrasp requires complete object point clouds, which is not the case for DexDiffuser nor FFHNet. The simulation results are presented in Table II. These results demonstrate that DexSampler, irrespective of the encoding method, consistently outperformed the FFHNet generator across all test datasets. The superior performance underscores the diffusion models’ effectiveness in synthesizing successful grasps. Moreover, DexSampler-BPS performed much better than DexSampler-PN2, indicating that the BPS encoding is less affected by point cloud irregularities, including missing parts and non-uniform density distribution. Finally, Fig. 2 shows some grasps generated by the DexSampler-BPS on the three datasets.
Apart from the success rate, we also report the grasp diversity and max penetration depth proposed in [4] in Table III. We excluded the Q1 grasp quality metric from [4] because, as reported in [4], that metric is unreliable when the max penetration depth exceeds 5mm, which is true for all our methods across all datasets. Following the definition in [4], the diversity metric is calculated as the average entropy across all joints. We add the diversity and max penetration depth of the training data generated using DexGraspNet as a reference. The results indicate that DexSampler outperforms both FFHNet and UniDexGrasp, highlighting its ability to generate diverse grasps with relatively low penetration depth.
Figure 2: Qualitative evaluation of DexSampler-BPS on objects from different datasets. The grasps, shown in pink, are generated on the partially observed point clouds, shown in red, rendered from the complete meshes, shown in blue.
@@ VI-B Real Robot Experiment
{adjustbox}
TABLE IV: Experimental results for real-world grasping. DexS represents DexSampler.
@@
Figure 3: Experimental objects. From 1 to 9: cracker box, sugar box, mustard bottle, bleach cleanser, sprayer, metal mug, goblet, toy plane, and pringles.
Figure 4: Successful and failed grasps from DexSampler-BPS-ESR-2. The first two images successfully grasp the bleach cleanser and the goblet, while the latter two show different failure cases on the metal mug and the pringles.
We evaluated the following methods: FFHNet, FFHNet-ESR-2, DexDiffuser-BPS, DexDiffuser-BPS-ESR-2, DexDiffuser-BPS-EGD and DexDiffuser-BPS-EGD-ESR-2. We omitted UniDexGrasp as it requires complete object point-clouds. It is worth highlighting that these methods are only trained on simulated data. The experimental results are summarized in Table IV. Like the simulation results, these results show that DexDiffuser consistently outperformed FFHNet across all test objects. The best method for picking real-world objects was DexSampler-BPS-ESR-2, and some of the grasps it produced are shown in Fig. 4. Interestingly, no methods successfully grasped the toy plane, mainly because the Allegro hand was not strong enough to keep the object secure when the last joint rotated $\pm$90°.
@@
We also report the inference time in Table IV. The iterative denoising process makes the diffusion-based methods slower at inference than FFHNet. For instance, the best performing DexDiffuser (DexS-BPS-ESR-2) is 3.5 times slower than the best performing FFHNet (FFHNet-ESR-2).